April 10, 2014

19:15 Pointing and collimation check

19:20 Telluric acquisition HIP35050

19:30 Struggling to find correct star because no finder were provided

19:50 Exposure for HIP35050

20:03 Acquisition of A611, the MOS unit failed

21:05 Problem with MOS. LUCI closed, switching to MODS

21:30 Moving to Greg's target A851 with MODS , seeing ~1", mask alignment

22:00 Started science exposure, avg seeing 0.95"

23:06 Acquisition CSW107

23:13 Some trouble with guide star, it is pretty faint for the weather, 15.7, seeing 0.96"

23:45 Acquisition CSWA107, changing the guide star

00:20 Lost the guide star with magnitude 14.9 due to clouds. Moving to CSWA104

01:12 moving to Greg's target MACSJ15224 for imaging , seeing ~1.4

01:53 Clouds: guide star lost

02:05 Exposure stopped due to clouds

03:12 trying MACSJ 15224 imaging, seeing ~0.9 . Still cloudy

03:58 Moving to CODEX 40313 imaging, seeing ~0.9. Clouds.

05:13 Taking calibrations

-- %USERSIG{GregoryLeeWalth - 2014-04-11}%

Comments


 
Topic revision: r4 - 14 Apr 2014, GregoryLeeWalth
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback